Find the local maximum and local minimum values of the function


f(x)= x^3-2x^2-4x+5

1
Expert's answer
2022-02-01T04:31:23-0500

Let us find the local maximum and local minimum values of the function "f(x)= x^3-2x^2-4x+5."

It follows that "f'(x)= 3x^2-4x-4=(x-2)(3x+2)." Therefore, "f'(x)=0" implies "x_1=2" and "x_2=-\frac{2}3." Taking into account that "f''(x)= 6x-4," we conclude that "f''(2)=8>0, \ f''(-\frac{2}3)=-8<0." Consequently, "x_1=2" is the point of a local minimum and "x_2=-\frac{2}3" is a the point of local maximum.